OpenStudy (anonymous):

HELP in 6th GRade Solve. a + 7 = 1 a =? A. 8 B. –8 C. –6 D. 6

OpenStudy (anonymous):

to get a, we have to substract 7 from both sides, like this a+7-7= 1-7

OpenStudy (anonymous):

i think the answer is C am i right?

OpenStudy (anonymous):

yes you are right :)

OpenStudy (charlotte123):

a + 7 = 1 - 7 - 7 a = 1 - 7 a = -6

OpenStudy (charlotte123):

So yes - You are correct! @angelv00 Great Job!

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

your welcome

OpenStudy (charlotte123):

No problem!